If P (1, 2), Q(4, 6), R(5, 7) and S(a, b) are the vertices of a parallelogram PQRS, then :
Options:
 

Grade:12th pass

1 Answers

Himanshu
7 years ago
Equate slope of PQ=RS & then by PS=QR.b-7/a-5=6-2/4-1. By solving this equation you get 3b=4a+1.And by equating slope of PS=QR b-2/a-1=7-6/5-4 . You get b=a+1.so by solving these 2 equations you get a=2&b=3.